# Parameters Inference for Nonlinear Wave Equations with Markovian Switching

### v1

@@ 5 Appendix
and both $\left\{\left(\theta^{\left(k_{s_{j}}\right)},\gamma^{\left(k_{s_{j}}\right)}\right)\right\}$ and $\left\{\left(\bar{\theta}^{\left(k_{s_{j}}+1\right)},\bar{\gamma}^{\left(k_{s_{j}}+1\right)}\right)\right\}$ are in $\operatorname{int}(\Omega)\cup\Omega_{2}$, they satisfy the following property: On $\left\{\bar{\gamma}_{i}^{(k)}\neq 0\right\}$, $\left|\bar{\theta}_{i}^{(k)}\right|/\bar{\gamma}_{i}^{(k)}$ is bounded. $\bar{\theta}_{i}^{(k)}$ and $\bar{\gamma}_{i}^{(k)}$ share the same zero properties.
@@
If $\left(\theta^{**},\gamma^{**}\right)\in\operatorname{int}(\Omega)$, then by the continuity of ^$\widehat{\mathcal{L}}\left(\theta,\gamma;\theta^{*},\gamma^{*}\right)$ on $\operatorname{int}(\Omega)$, the following holds:
@@
Choose a sufficiently small positive number $\varepsilon_{1}$ such that:
@@
According to equation 5.30, there exists a positive number $k_{1}$ such that for any $k>k_{1}$:
@@
Given the definition of ^$\widehat{\mathcal{L}}(\tilde{\theta},\tilde{\gamma};\theta,\gamma)$, we know that for $(\theta,\gamma)$, ^$\widehat{\mathcal{L}}(\tilde{\theta},\tilde{\gamma};\theta,\gamma)$ is continuous on $\Omega$. Therefore, there exists a positive number $k_{2}$ such that for any $k>k_{2}$:
@@
Combining the above parts, we can deduce that for any $k>\max\left\{k_{1},k_{2}\right\}$, the following holds:
@@
However, since:
$\left(\bar{\theta}^{(k)},\bar{\gamma}^{(k)}\right)\in\mathcal{G}\left(\theta^{(k)},\gamma^{(k)}\right),$ (5.36)
we know that:
@@
holds for any $(\theta,\gamma)\in\Omega$, leading to a contradiction with equation 5.35. Thus, we have:
@@
By equation 5.19, $\mathcal{G}$ is monotonically non-increasing with respect to $\mathcal{L}$. If $(\theta,\gamma)\notin S$, then $\forall(\bar{\theta},\bar{\gamma})\in\mathcal{G}$, $\mathcal{L}(\bar{\theta},\bar{\gamma})<\mathcal{L}(\theta,\gamma)$. In fact, if $\mathcal{L}(\bar{\theta},\bar{\gamma})=\mathcal{L}(\theta,\gamma)$, by equation 5.19, we know that ^$\widehat{\mathcal{L}}(\bar{\theta},\bar{\gamma};\theta,\gamma)=\widehat{\mathcal{L}}(\theta,\gamma;\theta,\gamma)$, and thus $(\theta,\gamma)\in S$. From equations 5.25 and 5.38, we have:
@@
According to the iteration generated by the mapping $\mathcal{G}$ in equation 5.3, there exists $z^{*}\in\partial\left|\theta^{*}\right|_{1}$ such that:
@@
$Proof:$ For the iterative sequence $\left\{\left(\theta^{(k)},\gamma^{(k)}\right)\right\}_{k=0}^{\infty}$ generated by the algorithm, one could verify that:
@@
$c^{(k)}$ $:=\operatorname{diag}\left[D^{T}\left(\sigma^{2}I_{n}+D\Gamma^{(k)}D^{T}\right)^{-1}D\right],$ (5.45)
$z^{(k+1)}$ $:=\left[z_{1}^{(k+1)},\cdots,z_{n}^{(k+1)}\right]^{T}\in\partial\left\|\theta^{(k+1)}\right\|_{1}.$ (5.46)
We introduce a function
@@
where $\tau$ is a fixed threshold. This function filters out the nonzero parameters smaller than $\tau$. When $\tau<\theta_{\text{min}}$, we have
$\theta_{\text{min}}:=\min_{i\in\mathbb{M}(\theta^{\text{true}})}|\theta_{i}^{\text{true}}|,$ (5.48)
and
$\mathbb{M}(\theta):=\{i:\theta_{i}^{\text{true}}\neq 0\}.$ (5.49)
Similar to Theorem 2.2, it is advantageous to consider error bound over the local time interval $[\tau_{k},\tau_{k+1})$. For $D^{T}D=I_{m}$, we have:
@@
According to Theorem 2.2, the sequence $\left\{\left(\theta^{(k)},\gamma^{(k)}\right)\right\}$ converges to the stationary point $\left(\theta^{*},\gamma^{*}\right)$ of 5.2. All stationary points of 5.2 satisfy the KKT conditions 5.40, hence $c^{(k)}$ converges to $c^{*}$ as $k\rightarrow\infty$. Taking the limit of 5.3 yields an implicit equation:
$\sigma^{2}c_{i}^{*}+\left|\theta_{i}^{*}\right|\sqrt{c_{i}^{*}}-1=0.$
Solving this quadratic equation for $\sqrt{c_{i}^{*}}$ gives:
@@
where $z^{*}:=\left[z_{1}^{*},z_{2}^{*},\cdots,z_{m}^{*}\right]^{T}$. Substituting 2.3 and 5.52 into 5.53, we get:
@@
For any $i\in\mathbb{M}\left(\theta^{\text{true}}\right)$, if $\left|D_{i}^{T}\varepsilon\right|<\frac{1}{2}\theta_{\min}$, then $i\in\mathbb{M}\left(\theta^{*}\right)$. Indeed, if $i\notin\mathbb{M}\left(\theta^{*}\right)$, then from 5.54, we have:
@@
$\left|\theta_{i}^{\text{real}}+D_{i}^{T}\varepsilon\right|\geq\left|\theta_{i}^{\text{real}}\right|-\left|D_{i}^{T}\varepsilon\right|\geq\frac{1}{2}\theta_{\min},$ (5.56)
@@
By assumption, this contradicts 5.55. For any $i\notin\mathbb{M}\left(\theta^{\text{real}}\right)$, if $\left|D_{i}^{T}\varepsilon\right|<\tau$, then $i\notin\mathbb{M}\left(\theta^{*}\right)$. Indeed, if $i\in\mathbb{M}\left(\theta^{*}\right)$, then from 5.54 and the assumption, we have:
@@
which contradicts the assumption. Thus, on $\left\{\max_{i}\left|D_{i}^{T}\varepsilon\right|<\tau\right\}$, we have $\mathbb{M}\left(\theta^{\text{true}}\right)=\mathbb{M}\left(\theta^{*}\right)$. Given $A^{T}A=I_{m}$, the values $A_{i}^{T}\varepsilon,i=1,\cdots,m$ are independent Gaussian random variables with mean zero and variance $\sigma^{2}$. By Hoeffding’s inequality, we have:
$\mathbb{P}\left[\left|D_{i}^{T}\varepsilon\right|\geq\xi\right]\leq 2e^{-\tau^{2}/4\sigma^{2}},$ (5.59)
and
$\mathbb{P}\left[\mathbb{M}\left(\theta^{*}\right)=\mathbb{M}\left(\theta^{\text{true}}\right)\right]$ $=\mathbb{P}\left[\max_{i}\left|D_{i}^{T}\varepsilon\right|<\tau\right]$
$=\prod_{i}\mathbb{P}\left[\left|D_{i}^{T}\varepsilon\right|<\tau\right]$
$\geq\left(1-2e^{-\tau^{2}/2\sigma^{2}}\right)^{m}$
$\geq 1-e^{-\tau^{2}/4\sigma^{2}}.$ (5.60)
For any $i\in\mathbb{M}\left(\theta^{\text{true}}\right)$, by Taylor’s theorem, there exists a real number $\eta\in(0,4\sigma^{2})$ such that:
@@
Substituting this into 5.54 gives:
@@
Finally, under the condition of Theorem 2.3, where $2\sigma\sqrt{\log 2m}\leq\tau<\frac{1}{2}\theta_{\min}-\sigma$, the uniform error bound is guaranteed. Observe that the RHS of 5.63 takes the form $x+\frac{a}{x}$, where $a>0$, indicating that $\tau$ attains its minimum within the first quadrant when $\tau=\sigma$. Assuming $2\sigma\sqrt{\log 2m}<\sigma$, it follows that $m\leq\frac{1}{2}e^{1/4}\approx 0.64$, which contradicts the fact that $m$ is a positive integer. Therefore, $\sigma<2\sigma\sqrt{\log 2m}<\frac{1}{2}\theta_{\min}-\sigma$. When the positive value $\tau$ is set to $\frac{1}{2}\theta_{\min}-\sigma$, there exists a consistent upper bound for the algorithm error. Substituting this into the error bound yields:

### v2

@@ 5 Appendix
and both $\left\{\left(\theta^{\left(k_{s_{j}}\right)},\gamma^{\left(k_{s_{j}}\right)}\right)\right\}$ and $\left\{\left(\bar{\theta}^{\left(k_{s_{j}}+1\right)},\bar{\gamma}^{\left(k_{s_{j}}+1\right)}\right)\right\}$ are in $\operatorname{int}(\Omega)\cup\Omega_{2}$. Then if $\left\{\bar{\gamma}_{i}^{(k)}\neq 0\right\}$, $\left|\bar{\theta}_{i}^{(k)}\right|/\bar{\gamma}_{i}^{(k)}$ is bounded. $\bar{\theta}_{i}^{(k)}$ and $\bar{\gamma}_{i}^{(k)}$ share the same zero properties.
@@
If $\left(\theta^{**},\gamma^{**}\right)\in\operatorname{int}(\Omega)$, then by the continuity of ^$\widehat{\mathcal{L}}\left(\theta,\gamma;\theta^{*},\gamma^{*}\right)$ on $\operatorname{int}(\Omega)$, the following holds
@@
Choose a sufficiently small positive number $\varepsilon_{1}$ such that
@@
According to equation 5.30, there exists a positive number $k_{1}$ such that for any $k>k_{1}$
@@
Given the definition of ^$\widehat{\mathcal{L}}(\tilde{\theta},\tilde{\gamma};\theta,\gamma)$, we know that for $(\theta,\gamma)$, ^$\widehat{\mathcal{L}}(\tilde{\theta},\tilde{\gamma};\theta,\gamma)$ is continuous on $\Omega$. Therefore, there exists a positive number $k_{2}$ such that for any $k>k_{2}$
@@
Combining the above parts, we can deduce that for any $k>\max\left\{k_{1},k_{2}\right\}$, the following holds
@@
However, since
$\left(\bar{\theta}^{(k)},\bar{\gamma}^{(k)}\right)\in\mathcal{G}\left(\theta^{(k)},\gamma^{(k)}\right).$ (5.36)
We have
@@
holds for any $(\theta,\gamma)\in\Omega$, leading to a contradiction with equation 5.35. Thus, we have
@@
By equation 5.19, $\mathcal{G}$ is monotonically non-increasing with respect to $\mathcal{L}$. If $(\theta,\gamma)\notin S$, then $\forall(\bar{\theta},\bar{\gamma})\in\mathcal{G}$, $\mathcal{L}(\bar{\theta},\bar{\gamma})<\mathcal{L}(\theta,\gamma)$. In fact, if $\mathcal{L}(\bar{\theta},\bar{\gamma})=\mathcal{L}(\theta,\gamma)$, by equation 5.19, we know that ^$\widehat{\mathcal{L}}(\bar{\theta},\bar{\gamma};\theta,\gamma)=\widehat{\mathcal{L}}(\theta,\gamma;\theta,\gamma)$, and thus $(\theta,\gamma)\in S$. From equations 5.25 and 5.38, we have
@@
According to the iteration generated by the mapping $\mathcal{G}$ in equation 5.4, there exists $z^{*}\in\partial\left|\theta^{*}\right|_{1}$ such that
@@
$Proof:$ For the iterative sequence $\left\{\left(\theta^{(k)},\gamma^{(k)}\right)\right\}_{k=0}^{\infty}$ generated by the algorithm, one could verify that
@@
$c^{(k)}$ $=\operatorname{diag}\left[D^{T}\left(\sigma^{2}I_{n}+D\Gamma^{(k)}D^{T}\right)^{-1}D\right],$ (5.45)
$z^{(k+1)}$ $=\left[z_{1}^{(k+1)},\cdots,z_{n}^{(k+1)}\right]^{T}\in\partial\left\|\theta^{(k+1)}\right\|_{1}.$ (5.46)
We define a truncation set
@@
where $\tau$ is a fixed threshold. When $\tau<\theta_{\text{min}}$, we have
$\theta_{\text{min}}=\min_{i\in\mathbb{M}(\theta^{\text{true}})}|\theta_{i}^{\text{true}}|.$ (5.48)
Similar to Theorem 2.2, it is advantageous to consider error bound over the local time interval $[\tau_{k},\tau_{k+1})$. For $D^{T}D=I_{m}$, we have
@@
According to Theorem 2.2, the sequence $\left\{\left(\theta^{(k)},\gamma^{(k)}\right)\right\}$ converges to the stationary point $\left(\theta^{*},\gamma^{*}\right)$ of 5.3. All stationary points of 5.3 satisfy the KKT conditions 5.40, hence $c^{(k)}$ converges to $c^{*}$ as $k\rightarrow\infty$. Taking the limit of 5.3 yields an implicit equation
$\sigma^{2}c_{i}^{*}+\left|\theta_{i}^{*}\right|\sqrt{c_{i}^{*}}-1=0.$ (5.51)
Solving this quadratic equation for $\sqrt{c_{i}^{*}}$ gives
@@
where $z^{*}:=\left[z_{1}^{*},z_{2}^{*},\cdots,z_{m}^{*}\right]^{T}$. Substituting 2.3 and 5.52 into 5.53, we get
@@
For any $i\in\mathbb{M}\left(\theta^{\text{true}}\right)$, if $\left|D_{i}^{T}\varepsilon\right|<\frac{1}{2}\theta_{\min}$, then $i\in\mathbb{M}\left(\theta^{*}\right)$. Indeed, if $i\notin\mathbb{M}\left(\theta^{*}\right)$, then from 5.54, we have
@@
$\left|\theta_{i}^{\text{true}}+D_{i}^{T}\varepsilon\right|\geq\left|\theta_{i}^{\text{true}}\right|-\left|D_{i}^{T}\varepsilon\right|\geq\frac{1}{2}\theta_{\min},$ (5.56)
and
@@
By assumption, this contradicts 5.55. For any $i\notin\mathbb{M}\left(\theta^{\text{true}}\right)$, if $\left|D_{i}^{T}\varepsilon\right|<\tau$, then $i\notin\mathbb{M}\left(\theta^{*}\right)$. Indeed, if $i\in\mathbb{M}\left(\theta^{*}\right)$, then from 5.54 and the assumption, we have
@@
which contradicts the assumption. Thus, on $\left\{\max_{i}\left|D_{i}^{T}\varepsilon\right|<\tau\right\}$, we have $\mathbb{M}\left(\theta^{\text{true}}\right)=\mathbb{M}\left(\theta^{*}\right)$. Given $D^{T}D=I_{m}$, the values $D_{i}^{T}\varepsilon,i=1,\cdots,m$ are independent Gaussian random variables with mean zero and variance $\sigma^{2}$. By Hoeffding’s inequality, we have
$\mathbb{P}\left[\mathbb{M}\left(\theta^{*}\right)=\mathbb{M}\left(\theta^{\text{true}}\right)\right]=\mathbb{P}\left[\max_{i}\left|D_{i}^{T}\varepsilon\right|<\tau\right]$ (5.59)
$=$ $\prod_{i}\mathbb{P}\left[\left|D_{i}^{T}\varepsilon\right|<\tau\right]\geq\left(1-2e^{-\tau^{2}/2\sigma^{2}}\right)^{m}\geq 1-e^{-\tau^{2}/4\sigma^{2}}.$
For any $i\in\mathbb{M}\left(\theta^{\text{true}}\right)$, by Taylor’s theorem, there exists $\eta\in(0,4\sigma^{2})$ such that
@@
Substituting this into 5.54 gives
@@
Finally, under the condition of Theorem 2.3, where $2\sigma\sqrt{\log 2m}\leq\tau<\frac{1}{2}\theta_{\min}-\sigma$, the uniform error bound is guaranteed. Observe that the RHS of 5.62 takes the form $x+\frac{a}{x}$, where $a>0$, indicating that $\tau$ attains its minimum within the first quadrant when $\tau=\sigma$. Assuming $2\sigma\sqrt{\log 2m}<\sigma$, it follows that $m\leq\frac{1}{2}e^{1/4}\approx 0.64$, which contradicts the fact that $m$ is a positive integer. Therefore, $\sigma<2\sigma\sqrt{\log 2m}<\frac{1}{2}\theta_{\min}-\sigma$. When the positive value $\tau$ is set to $\frac{1}{2}\theta_{\min}-\sigma$, there exists a consistent upper bound for the algorithm error. Substituting this into the error bound we yield that